When the Cu2+ concentration is 6.58×10-4 M, the observed cell potential at 298K for an electrochemical cell with the following reaction is 0.979V. What is the Cr3+ concentration?

3Cu2+(aq) + 2Cr(s)3Cu(s) + 2Cr3+(aq)

Answer:  M

When the Cu2+ concentration is 2.71×10-4 M, the observed cell potential at 298K for an electrochemical cell with the following reaction is 0.971V. What is the Cr3+ concentration?

3Cu2+(aq) + 2Cr(s)3Cu(s) + 2Cr3+(aq)

Answer:  M
Answer #1

3 lut + 2 Cr 3 cu 2+ 2 en 3+ Standarda reduction of Potential of half cens: Elu24/ eu = 0.34 v -0.74 V Eleretler Eeen= E euroGiven, E [eu 2+ = 6.58x10 PM and Ereu 0.979 v. [724) 0.979 1.08 0.05 log 2 [er3t] = 1.08 – 0.059 (3-5) (6.58x169) log [2.83+7L Cr3+) = 1.56 Concentration, of (20 37 1.56/m